9) A rectangular channel 8 feet wide carries 2.5 ft/s of water at a depth of 3.2 feet.

a) What are the most efficient attributes (d, A, P, and R) of the channel?

b) If a 1-foot wide pier is placed in the middle of the channel, find the elevation of the water surface at the pier. What is the minimum width of the constriction that will not cause a rise in the upstream water surface?
Answer #1

Here for the last part of 2nd part one concept is used that is specific energy is constant for both section .E =d +v^2/2g

For rectangular channel Ec=1.5*dc

dc=(q^2/g)^1/3

Comment for any doubt about this solution thank you

Don't forget to upvote if you got your answer